|  October 30, 2014  |  
Overcast   24.0F  |  Forecast »
Bookmark and Share Email this page Email Print this page Print Feed Feed

Repsol Q2 Pad Gas and Mud Release SITREP # 19

ALASKA DEPARTMENT OF ENVIRONMENTAL CONSERVATION
Division of Spill Prevention and Response
Prevention and Emergency Response Program

SITUATION REPORT  SPILL NUMBER: 12399904601
TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 a.m. on February 15, 2012.

 

Repsol Q2 Pad Gas and Mud Release  SITREP # 19
TIME/DATE OF SITUATION REPORT: 3:30 p.m. on March 3, 2012

RESPONSE ACTION: Temperatures remain extremely cold. With an ambient temperature of -44°F and wind chill of -68°F, use of equipment outside the drilling rig continues to be limited.

The first well log, made last night, showed a blockage of the drill pipe. A coiled tubing unit and crane have been contracted to clean out the drill pipe, but the temperature must be above -35°F for them to operate. The weather is not forecasted to warm above -35°F until this Tuesday, March 6. The contracted coiled tubing unit is currently working on another job, which will take about one day to complete once the temperatures warm up above -35°F.

Once the coiled tubing unit completes its present job, it will be moved to the Q2 pad, rigged and used to clean out the blockage. Once the blockage is cleared, the wireline crew will run another well logging. If no further problems are discovered, the information gained in the second well logging will be used to finalize the well-kill plan.

Repsol reports that 68 barrels (2,856 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,800 barrels (75,600 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid collected is condensed water from the steam used to thaw and clean the rig. Another 178 cubic yards of solid waste (frozen drilling mud) were manifested and shipped offsite yesterday, bringing the total to 842 cubic yards of solid waste that have been manifested and shipped offsite.

The Unified Command decided to temporarily close the Anchorage Incident Command Post (ICP) until Tuesday morning, March 6, 2012, due to the temperature hold on the wellbore clean out operation. ADEC responders at Palm Pad will remain in place, and ADEC responders working in Anchorage will work out of the local ADEC office until the ICP reopens. If the temperature moderates earlier than currently predicted, the ICP will reopen earlier as well.

WEATHER: Today: Partly cloudy with a few flurries. Patchy ice fog. Highs around 35 below. Southwest winds around 10 mph.
Tonight: Partly cloudy. Areas of ice fog. Lows 40 to 45 below. Southwest winds around 10 mph.

TIME/DATE OF THE NEXT SITUATION REPORT: 3:30 p.m. on March 06, 2012

 

SITREP # 18
SPILL NUMBER: 12399904601
TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 AM on February 15, 2012.
TIME/DATE OF SITUATION REPORT: 3:30 PM on March 2, 2012

RESPONSE ACTION: Response crews continue 24-hour operations to clean and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

The Unified Command (UC) traveled to Nuiqsut for a public meeting yesterday, March 2, 2012. About 30 people attended the meeting at which the UC gave information and answered questions about what had occurred, and current and planned operations.

Temperatures remain extremely cold with a -63°F wind chill, so outside equipment use continues to be limited. However, work inside the rig continues.

Crews have completed the cleaning and drying of the rig’s power components and the rig is operational.

A wireline is used to lower instruments down the borehole of the well to gather information on the current condition of the well; a process called well logging. The wireline crew has initiated this process. The information gathered in the well logging will be used for finalization of the well-kill plan.

Repsol reports that 8 barrels (336 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,732 barrels (72,744 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid collected is condensed water from the steam used to thaw and clean the rig. Another 114 cubic-yards of solid waste (frozen drilling mud) were manifested and shipped offsite yesterday bringing the total to 664 cubic-yards of solid waste that have been manifested and shipped offsite.

WEATHER: Today: Mostly sunny. Patchy ice fog. Highs 25 to 35 below. Variable winds less than 10 mph.
Tonight: Partly Cloudy. Patchy ice fog. Lows 40 below to 50 below. Light winds.

TIME/DATE OF THE NEXT SITUATION REPORT: 3:30 PM on March 03, 2012

 

SITREP # 17
SPILL NUMBER: 12399904601
TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 AM on February 15, 2012.
TIME/DATE OF SITUATION REPORT: 3:30 PM on March 1, 2012

RESPONSE ACTION: Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

Temperatures remain extremely cold with a -58°F wind chill, so outside equipment continues to be limited. However, work inside the rig continues.

Crews are continuing the cleaning and drying of the rig’s power components; critical steps in the eventual re-start of the rig.
The re-start of the rig is one in a series of steps in the progression needed for completion of the well-kill operation. However, due to the need for a very methodical and careful cleaning and drying of the rig’s power components, the time of the re-start cannot be determined at this time.

Work on setting up the wireline equipment for well inspection has begun. The wireline is used to lower instruments down the borehole of the well to gather information on the current condition of the well. The information gathered will be used for finalization of the well-kill plan.
Repsol reports that 58 barrels (2,436 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day,

bringing the total to 1,724 barrels (72,408 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. Another 75 cubic yards of solid waste (frozen drilling mud) were also manifested and shipped offsite yesterday bringing the total to 550 cubic yards of solid waste that have been manifested and shipped offsite.

WEATHER: Today: Mostly sunny. Patchy ice fog. Highs 30 to 35 below. East winds around 10 mph.
Tonight: Mostly clear. Patchy ice fog. Lows 35 to 45 below. Northeast winds around 10 mph.

TIME/DATE OF THE NEXT SITUATION REPORT: 3:30 PM on March 02, 2012

For complete and ongoing information, go to ADEC website

Add your comment:
Advertisement
Advertisement
Advertisement
Advertisement
Advertisement